OpenStudy (anonymous):

Find the value of X: 4x^2-44=5x

OpenStudy (zepp):

\(4x^2-5x-44=0\) Send 5x to the other side. Then use the quadratic formula to find the value of x.

OpenStudy (anonymous):

Thats what i need help on.. whats the quadratic answer.... ( ) ( )

OpenStudy (anonymous):

its 5-5\[(5- or +5\sqrt{11})\div8\]

OpenStudy (anonymous):

sorry the second part

OpenStudy (anonymous):

no what i got was (4x+11) (x-4) Which would become -16x+11x= -5x So then i got -2.75, 4 as the values of x..